OpenStudy (alyssa_marie_gunn1):

Which products result in a difference of squares? Check all that apply. (5z + 3)(–5z – 3) (w – 2.5)(w + 2.5) (8g + 1)(8g + 1) (–4v – 9)(–4v + 9) (6y + 7)(7y – 6) (p – 5)(p – 5) @aveline

jimthompson5910 (jim_thompson5910):

Difference of squares rule \[\Large a^2 - b^2 = (a+b)(a-b)\]

jimthompson5910 (jim_thompson5910):

Example: \[\Large x^2-25 = x^2-5^2 = (x-5)(x+5)\]

OpenStudy (ciarabelen):

I believe that only numbers 2 and 4 are those that are correct for this question.